# Comparative Performance of Collaborative Bandit Algorithms: Effect of Sparsity and Exploration Intensity

### v1

@@ Abstract
This paper offers a comprehensive analysis of collaborative bandit algorithms and provides a thorough comparison of their performance. Collaborative bandits aim to improve the performance of contextual bandits by introducing relationships between arms (or items), allowing effective propagation of information. Collaboration among arms allows the feedback obtained through a single user (item) to be shared across related users (items). Introducing collaboration also alleviates the cold user (item) problem, i.e., lack of historical information when a new user (item) arriving to the platform with no prior record of interactions. In the context of modeling the relationships between arms (items), there are two main approaches: Hard and soft clustering. We call approaches that model the relationship between arms in an absolute manner as hard clustering, i.e., the relationship is binary. Soft clustering relaxes membership constraints, allowing fuzzy assignment. Focusing on the latter, we provide extensive experiments on the state-of-the-art collaborative contextual bandit algorithms and investigate the effect of sparsity and how the exploration intensity acts as a correction mechanism. Our numerical experiments demonstrate that controlling for sparsity in collaboration improves data efficiency and performance as it better informs learning. Meanwhile, increasing the exploration intensity acts as a correction because it effectively reduces variance due to potentially misspecified relationships among users. We observe that this misspecification is further remedied by introducing latent factors, and thus, increasing the dimensionality of the bandit parameters.
@@
Recommender Systems, Contextual Bandits, Reinforcement Learning, Multi-armed Bandits, Online Recommendation, Offline Evaluation Methods.
@@ 1 Introduction
Personalizing user experience on recommender platforms has seen growing interest from various fields, such as graph neural networks (Yin et al., , 2019; Wu et al., , 2022; Gao et al., , 2023; Sharma et al., , 2024), reinforcement learning (Zou et al., , 2019; Xin et al., , 2020; Afsar et al., , 2022; Wang et al., , 2024), and more recently, generative models (Liang et al., , 2018; Wu et al., , 2019; Sachdeva et al., , 2019; Deldjoo et al., , 2024), with many also focusing on their practical applications (Bobadilla et al., , 2013; Lu et al., , 2015; Guo et al., , 2022). The personalization task requires an online platform to effectively utilize the information it can collect on users’ interactions with the platform (e.g., using their session logs) to generate insights about their preferences. As new information on users may emerge, or user preferences may shift, the platform needs an algorithmic approach that can quickly and reliably process new information and revise strategies accordingly. Hence, most platforms interacting with users implement online learning solutions to navigate this space effectively (Zhou et al., , 2020). The effectiveness of such systems rely on finding high-quality mappings (embeddings) between the users and the recommended items.
@@
The main objective of these platforms is to ensure that users stay engaged and remain on the platform for longer, enabling higher value generation. Therefore, recommendations should raise users’ engagement, while allowing the platform to collect new information about users’ preferences. Such information will enable the platform to improve its recommender system strategies. This arises the so-called exploration-exploitation trade-off commonly associated with the multi-armed bandits problem in the literature. While the platform tends to optimize the information it already has about users to recommend them the best items (exploit), it also needs to capture information about users with shifting interests or emerging items (explore) to ensure that users are kept engaged. For example, exploration in an online advertisement setting, where ads are to be shown to users, may mean that the platform should devote a certain portion of its efforts on recommending less relevant ads to users to learn more about their preferences. In the remainder of this paper, we will keep this online advertisement setting as our running example.
@@
In this work, we focus on analyzing this exploration-exploitation trade-off using multi-armed bandit algorithms, which provide a principled solution for this dilemma. Formally, we say that an arm is an item that can be recommended (assigned) to a user. The goal of the decision-maker (learner) is to recommend specific items to specific users in order to maximize the instances of users choosing to interact with the items. We assume that there is a set of items and a static set of users with (deterministic) information known to the learner at the time of interaction and an unknown parameterized function (usually with a parameter vector $\bm{\theta}$) that maps the best item to a specific user. The decision-maker aims to learn this function by continuously interacting with users over the course of a fixed and known time horizon, and to minimize the regret, which is the cost incurred due to uncertainty. The LinUCB algorithm due to Li et al., (2010) provides a precise approach for this setting, and therefore we use (a version of) this algorithm as a baseline to evaluate the value of more complicated algorithms.
@@
Starting with the above setting of utilizing contextual information, one may imagine numerous models here: (1) use contextual information on items and associate users with bandit parameters, (2) use contextual information on users and associate items with bandit parameters, (3) combine contextual information of items and users (i.e., interactions of different contexts), and associate items or users with bandit parameters. [^1] An algorithm for the setting of combining contextual information of items and users is presented in Li et al., (2010) as the hybrid model. We assume that there are many more users than items in our system. Hence, we seek models that relate different users’ behaviors to each other, which helps to (1) reduce the amount of information kept in and fed to the model, and, (2) remedy the cold start problem, i.e., users with no past history of interactions in the system. Through this relation, a new user’s preferences can be modeled using the preferences of existing users.
In online advertisements, the reward signal, i.e., if a user clicked on an ad or not, is both sparse and has high variance. Former is because many of the users do not click on ads, and latter is because among the ones who click, a considerable portion click on accident and not necessarily intend to interact with the ad itself (Thomas et al., , 2015). Hence, it is important to identify unique users, both to track how they evolve as time advances and to better capture their underlying preferences, which helps in establishing long-term relationships with the users, maximizing the payoff. In this example, we set payoff to $1$ if the user clicks on the shown ad, and 00 otherwise.
Note that, even though the users may be identified (which is getting harder due to privacy protections and laws, e.g., GDPR Cookie Law222gdpr.eu/cookies), the signals may still be sparse, and, a policy interacting with new users may be inefficient if they primarily rely on user features.
For reasons mentioned above, we focus on modeling the relationship between different users to both increase the data efficiency, potentially reducing learning complexity and to facilitate the problem of cold users. The idea of collaboration may also improve performance with new items recommendation, and a similar approach may be utilized to map items to items to solve the cold item problem.
@@
In order to confirm that introducing a collaboration structure indeed helps algorithms to be more data efficient, and therefore improve their performance, we conduct comprehensive experiments to compare and contrast M-LinUCB with the CoLin (Wu et al., 2016a, ) and FactorUCB (Wang et al., , 2017) algorithms. [^3] M-LinUCB is a version of the LinUCB algorithm that is more suitable for our experiments (Wu et al., 2016a, ). CoLin and FactorUCB algorithms model the relationships among users through a similarity matrix. In our experiments, we demonstrate that controlling for sparsity in this user-similarity matrix further improves data efficiency and performance as it informs relationships among users more effectively. Additionally, FactorUCB introduces latent factors to expand the dimensionality of the item feature space. We observe that the introduction of these latent factor helps FactorUCB outperform CoLin by a large margin, and we argue that this improvement is due to potential misspecification of the collaboration structure, i.e., the similarity matrix used in experiments. In order to test the effects of this potential misspecification further, we investigate the exploration intensity, and analyze the extent to which it acts as a correction. We argue that increasing the exploration intensity helps as it balances the variance due to misspecification through putting more emphasis on exploratory actions. Note that this behavior may not be consistent across different paradigms where underlying distributions quickly stabilize.
@@
We use the Yahoo! Today Module (2009) dataset (version R6A) in our experiments (Li et al., , 2010). [^4] Obtained from webscope.sandbox.yahoo.com/catalog.php?datatype=r This dataset is considerably large with over $45$ million data points, $20$ candidate articles, and a total of $270$ articles over the course of a ten-day period. The payoff of the interaction of the user with the recommended article is called a click, as described earlier, and is $1$ if the user clicks on the shown article, and 00 otherwise. Also note that, because this dataset is de-identified, we cannot treat users independently and we cluster users before running any of the algorithms we analyze.
@@
The rest of the paper is organized as follows: In Section 2, we survey the relevant works. In Section 3, we present the preliminaries and the algorithms we use in our experiments. We extensively analyze the results in Section 4 and set up for Section 5 where we discuss our findings and proposed future directions, both empirical and technical, in detail.
2 Related Literature
We examine two principal types of algorithms in this field, namely offline and online learning algorithms. We categorize collaborative filtering and matrix factorization based algorithms as offline (we refer the reader to some online applications of these algorithms (Zangerle and Bauer, , 2022; Raza and Ding, , 2022; Chen et al., , 2023)). Then, we survey the collaboration-enabled algorithms under online learning. This collaboration is generally done through (weighted) graphs. Here, we consider clustering as a graph-based model.
2.1 The Offline Setting: Collaborative Filtering and Matrix Factorization
Collaborative Filtering (CF) (Goldberg et al., , 1992) is one of the more popular approaches in recommender systems, mostly due to its data efficiency as it either uses (1) a neighborhood based model to find relationships between items or between users and make recommendations based on this shared knowledge graph to relate items to items, or, (2) a latent factor model, e.g., Matrix Factorization (MF), to embed items and users to the same latent factor space and compare both in that space.
@@ 2.1 The Offline Setting: Collaborative Filtering and Matrix Factorization
2.2 The Online Setting: Collaboration in Bandits
@@
We focus on three main works: LinUCB algorithm (Li et al., , 2010), CoLin algorithm (Wu et al., 2016a, ), and FactorUCB algorithm (Wang et al., , 2017). The LinUCB algorithm introduces the use of contextual bandit modeling. CoLin builds onto LinUCB by introducing user neighborhoods to relate users to each others. Furthermore, FactorUCB builds onto CoLin by incorporating latent factors. In our experiments, we analyze and comment on this evolution of models, providing additional context on each of these algorithms and how they achieve the observed improvements.
@@ 3 Problem Formulation
We first review fundamentals of contextual bandit algorithms. We describe the preliminaries on contextual bandits and set the notation for the remainder of the paper. Next, we introduce the collaboration structure into contextual bandit algorithms and present the three algorithms relevant to this work, namely LinUCB (Li et al., , 2010), CoLin (Wu et al., 2016a, ), and FactorUCB (Wang et al., , 2017).
@@
At each time $t\in[T]$, a user $u_{t}$ arrives to the system with context $u_{t}(x)\in\mathbb{R}^{d_{u}}$.555$[n]$ denotes the natural numbers up to and including $n$, i.e., $[n]=\{1,2,\dots,n\}$. Context of each arm in $\mathcal{A}$ at time $t$, $a_{t}(x)\in\mathbb{R}^{d_{a}}$, may be combined with context of user $u_{t}$ to obtain $\bm{x}_{a_{t},u_{t}}\in\mathbb{R}^{d}$, where $d=d_{u}+d_{a}$. [^6] Note that the context of the user may change across time, e.g., number of recommendations they received so far may be one of the information dimensions (Theocharous et al., , 2015). We assume that $d_{u}$ and $d_{a}$ remain unchanged throughout the time horizon. After the decision-maker (learner) pulls arm $a_{t}$ at time $t$, they receive a payoff of $r_{a_{t},u_{t}}$ from user $u_{t}$. This payoff is assumed to be a function of the context and the system state at time $t$, i.e., $r_{a_{t},u_{t}}=f(\bm{x}_{a_{t},u_{t}},\bm{S}_{t})$, where $\bm{S}_{t}$ may contain any information collected up to time $t$, and $f$ is an arbitrary function. Note that the learner has access to $\bm{S}_{0}$ prior to any interactions.
@@
Contextual bandit models provide a framework to incorporate the abundant information collected from users and items (i.e., arms) into decision making. In terms of interacting with users, by taking an action (such as recommending an item) and capturing the user’s feedback, most algorithms work in a similar manner. The distinguishing factor across them is how they utilize the collected information and propagate it across all users and actions. To materialize this, we introduce Algorithm 1 as a basis for all contextual bandit algorithms.
1 Input: Set of parameters $\mathcal{C}$, set of arms $\mathcal{A}_{t}$ for $t\in[T]$, initial system information $\bm{S}_{0}$, function $g(\cdot)$
@@ 3.1 Contextual bandits in collaboration
1 Input: $\alpha\in\mathbb{R}_{+}$, initial arm set $\mathcal{A}_{0}$
@@ 16       $\mathbf{b}_{a_{t}}\leftarrow\mathbf{b}_{a_{t}}+r_{a_{t},u_{t}}\bm{x}_{a_{t},u_{t}}$
1 Input: $\alpha\in\mathbb{R}_{+}$, initial item set $\mathcal{A}_{0}$, user cluster set $\mathrm{C}\triangleq\cup_{i=1}^{M}\mathrm{C}_{i}$.
@@ 15      Pull arm $a_{t}=\arg\max_{a\in\mathcal{A}_{t}}p_{t,a}$ and observe reward $r_{a_{t},u_{t}}$
Notice that we associate arms with bandit parameters in Algorithm 2. To extend this to a setting where arms are longer-lived than users, we instead associate users with bandit parameters and use the contextual information on arms to design the best policy. Additionally, we consider a setting where there are prohibitively many users and we remedy this by clustering users and associating these clusters with bandit parameters. In this setting, whenever a user belonging to a specific cluster gets an arm assigned to them, it is as if all the other users in the same cluster do so. This also helps in propagating the information collected on a user to another one based on their perceived similarity. Assuming that a structured way to extract such similarity information exists, such as already existing friend networks in social media, collaboration propagates new information and thus speeds up the learning process, and also offers a remedy to the cold-start problem in recommender systems. The intuition is similar to that in Li et al., (2010): The payoff from a user’s interaction may be a function of both their own preference and the influence they experience from others.
There are two main approaches in the literature to inject collaboration into contextual bandit models, namely hard- and soft-clustering. In hard-clustering, connected arms are assumed to have similar or exactly the same bandit parameters. For instance, GOB.Lin (Cesa-Bianchi et al., , 2013) requires connected users in a network to share similar bandit parameters through a graph Laplacian-based regularization, i.e., the bandit parameters across connected users must be close to each other in space. CLUB (Gentile et al., , 2014) makes a similar but stronger modeling assumption and forces arms in the same cluster to have exactly the same bandit parameters. In soft-clustering, arms are connected through a weighted adjacency graph, where nodes are contextual bandits on users and the weights on the edges govern the influence between pairs of users. In upcoming parts, we introduce the notion of sparsity in these settings, and say that we have $100\%$ sparsity if the weighted adjacency graph keeps all weights non-zero. A $50\%$ sparsity then means that the weighted adjacency graph keeps only the largest half of the weights non-zero. [^9] This is from each user’s perspective, i.e., this is a directed graph and not necessarily symmetric in both directions between nodes.
1 Input: $\alpha\in\mathbb{R}_{+}$, $\mathcal{A}_{0}$, $\mathbf{W}\in\mathbb{R}^{M\times M}$.

### v2

@@ Abstract
This paper offers a comprehensive analysis of collaborative bandit algorithms and provides a thorough comparison of their performance. Collaborative bandits aim to improve the performance of contextual bandits by introducing relationships between arms (or items), allowing effective propagation of information. Collaboration among arms allows the feedback obtained through a single user (item) to be shared across related users (items). Introducing collaboration also alleviates the cold start problem, i.e., lack of historical information when a new user arriving to the platform with no prior record of interactions. In the context of modeling the relationships between arms, there are two main approaches: Hard and soft clustering. We call approaches that model the relationship between arms in an absolute manner as hard clustering, i.e., the relationship is binary. Soft clustering relaxes membership constraints, allowing fuzzy assignment. Focusing on the latter, we provide extensive experiments on the state-of-the-art collaborative contextual bandit algorithms and investigate the effect of sparsity and how the exploration intensity acts as a correction mechanism. Our numerical experiments demonstrate that controlling for sparsity in collaboration improves data efficiency and performance as it better informs learning. Meanwhile, increasing the exploration intensity acts as a correction because it effectively reduces variance due to potentially misspecified relationships among users. We observe that this misspecification is further remedied by introducing latent factors, and thus, increasing the dimensionality of the bandit parameters.
@@
Recommender Systems, Contextual Bandits, Reinforcement Learning, Multi-armed Bandits.
@@ 1 Introduction
Personalizing user experience on recommender platforms has seen growing interest from various fields, such as graph neural networks (Yin et al., , 2019; Wu et al., , 2022; Gao et al., , 2023; Sharma et al., , 2024), reinforcement learning (Zou et al., , 2019; Xin et al., , 2020; Afsar et al., , 2022; Wang et al., , 2024), and more recently, generative models (Liang et al., , 2018; Wu et al., , 2019; Sachdeva et al., , 2019; Deldjoo et al., , 2024), with many also focusing on their practical applications (Bobadilla et al., , 2013; Lu et al., , 2015; Guo et al., , 2022). The personalization task requires an online platform to effectively utilize the information it can collect on users’ interactions with the platform (e.g., using their session logs) to generate insights about their preferences. As new information on users may emerge, or user preferences may shift, the platform needs an algorithmic approach that can quickly and reliably process new information and revise strategies accordingly. Hence, most platforms interacting with users implement online learning solutions to navigate this space effectively (Zhou et al., , 2020). The effectiveness of such systems relies on finding high-quality mappings (embeddings) between the users and the recommended items.
@@
The main objective of these platforms is to ensure that users stay engaged and remain on the platform for longer, enabling higher value generation. Therefore, recommendations should raise users’ engagement, while allowing the platform to collect new information about users’ preferences. This data helps the platform refine its recommender strategies. This gives rise to the classic exploration-exploitation trade-off often associated with the multi-armed bandit problem in the literature. While the platform tends to optimize the information it already has about users to recommend them the best items (exploit), it also needs to capture information about users with shifting interests or emerging items (explore) to ensure that users are kept engaged. For example, exploration in an online advertisement setting, where ads are to be shown to users, may mean that the platform should devote a certain portion of its efforts on recommending potentially less relevant ads to users to learn more about their preferences. This online advertisement setting will serve as our running example for the rest of the paper.
@@
In this work, we focus on analyzing the exploration-exploitation trade-off using multi-armed bandit algorithms, which provide a principled solution for this dilemma. Formally, we say that an arm is an item that can be recommended (assigned) to a user. The goal of the decision-maker (learner) is to recommend specific items to specific users in order to maximize the instances of users choosing to interact with the items. We assume a set of items and a static set of users with (deterministic) information known to the learner at the time of interaction and an unknown parameterized function (usually with a parameter vector $\bm{\theta}$) that maps the best item to a specific user. The decision-maker aims to learn this function by continuously interacting with users over the course of a fixed and known time horizon, and to minimize the regret, which is the cost incurred due to uncertainty. Using a version of the LinUCB algorithm Li et al., (2010) as a baseline, we precisely evaluate more complex methods for this problem.
@@
Starting with the above setting of utilizing contextual information, one may imagine numerous models here: (1) use contextual information on items and associate users with bandit parameters, (2) use contextual information on users and associate items with bandit parameters, (3) combine contextual information of items and users (i.e., interactions of different contexts), and associate items or users with bandit parameters. [^1] An algorithm for the setting of combining contextual information of items and users is presented in Li et al., (2010) as the hybrid model. We assume that there are many more users than items in our system. Hence, we seek models that relate different users’ behaviors to each other, which helps to (1) reduce the amount of information kept in and fed to the model, and, (2) remedy the cold-start problem, i.e., users with no past history of interactions in the system. Through this relation, a new user’s preferences can be modeled using the preferences of existing users.
In online advertising, the reward signal—whether or not a user clicked on an ad—is both sparse and highly variable. It is sparse because few users click on ads, and it has high variance because many of those clicks are accidental, not intentional (Thomas et al., , 2015). Therefore, identifying unique users is crucial. This allows us to track how their preferences evolve over time, capture their underlying interests more accurately, and ultimately build long-term relationships that maximize payoff. In this model, we define the payoff as 1 if a user clicks on an ad, and 0 otherwise. Note that, even though the users may be identified (which is getting harder due to privacy protections and laws, e.g., GDPR Cookie Law)222gdpr.eu/cookies, the signals may still be sparse, and a policy interacting with new users may be inefficient if it primarily relies on user features. For reasons mentioned above, we focus on modeling the relationship between different users to both increase the data efficiency, potentially reducing learning complexity and to facilitate the problem of cold users. The idea of collaboration may also improve performance with new items recommendation, and a similar approach may be utilized to map items to items to solve the cold item problem.
@@
In order to confirm that introducing a collaboration structure indeed helps algorithms to be more data efficient, and therefore improve their performance, we conduct comprehensive experiments to compare and contrast M-LinUCB with the CoLin (Wu et al., 2016a, ) and FactorUCB (Wang et al., , 2017) algorithms. [^3] M-LinUCB is a version of the LinUCB algorithm that is more suitable for our experiments (Wu et al., 2016a, ). CoLin and FactorUCB algorithms model the relationships among users through a similarity matrix. In our experiments, we demonstrate that controlling for sparsity in this user-similarity matrix further improves data efficiency and performance as it informs relationships among users more effectively. Additionally, FactorUCB introduces latent factors to expand the dimensionality of the item feature space. We observed that adding latent factors helped FactorUCB significantly outperform CoLin. We believe this is because the latent factors compensate for a potentially inaccurate collaboration structure, which is represented by the similarity matrix. In order to test the effects of this potential misspecification further, we investigate the exploration intensity, and analyze the extent to which it acts as a correction. We argue that increasing the exploration intensity helps as it balances the variance due to misspecification through putting more emphasis on exploratory actions. Note that this behavior may not be consistent across different paradigms where underlying distributions quickly stabilize.
@@
We use the Yahoo! Today Module (2009) dataset (version R6A) in our experiments (Li et al., , 2010). [^4] Obtained from webscope.sandbox.yahoo.com/catalog.php?datatype=r This dataset is considerably large with over $45$ million data points, $20$ candidate articles, and a total of $270$ articles over the course of a ten-day period. The payoff of the interaction of the user with the recommended article is called a click, as described earlier, and is $1$ if the user clicks on the shown article, and 0 otherwise. Also note that, because this dataset is de-identified, we cannot treat users independently and we cluster users before running any of the algorithms we analyze.
@@
The rest of the paper is organized as follows: In Section 2, we survey the relevant works. In Section 3, we present the preliminaries and the algorithms we use in our experiments. We extensively analyze the results in Section 4 and Section 5 discusses our findings and proposed future directions, both empirical and technical, in detail.
2 Related Work
We examine two principal types of algorithms in this field, namely offline and online learning algorithms. We categorize collaborative filtering and matrix factorization based algorithms as offline (though we refer the reader to some online applications of these algorithms (Zangerle and Bauer, , 2022; Raza and Ding, , 2022; Chen et al., , 2023)). Then, we survey the collaboration-enabled algorithms under online learning. This collaboration is generally done through (weighted) graphs. Here, we consider clustering as a graph-based model.
2.1 Offline Setting: Collaborative Filtering and Matrix Factorization
Collaborative Filtering (CF) (Goldberg et al., , 1992) is one of the more popular approaches in recommender systems, mostly due to its data efficiency. CF either uses (1) a neighborhood-based model to find relationships between items or between users and make recommendations based on this shared knowledge graph to relate items to items, or, (2) a latent factor model, e.g., Matrix Factorization (MF), to embed items and users to the same latent factor space and compare both in that space.
@@ 2.1 Offline Setting: Collaborative Filtering and Matrix Factorization
2.2 Online Setting: Collaboration in Bandits
@@
We focus on three main works: LinUCB algorithm (Li et al., , 2010), CoLin algorithm (Wu et al., 2016a, ), and FactorUCB algorithm (Wang et al., , 2017). The LinUCB algorithm introduces the use of contextual bandit modeling. CoLin builds onto LinUCB by introducing user neighborhoods to relate users to each other. Furthermore, FactorUCB builds onto CoLin by incorporating latent factors. In our experiments, we analyze and comment on this evolution of models, providing additional context on each of these algorithms and how they achieve the observed improvements.
@@ 3 Problem Formulation
We first review fundamentals of contextual bandit algorithms. We describe the preliminaries on contextual bandits and set the notation for the remainder of the paper. Next, we introduce the collaboration structure into contextual bandit algorithms and present the three algorithms relevant to this work: LinUCB (Li et al., , 2010), CoLin (Wu et al., 2016a, ), and FactorUCB (Wang et al., , 2017).
@@
At each time $t\in[T]$, a user $u_{t}$ arrives to the system with context $u_{t}(x)\in\mathbb{R}^{d_{u}}$.555[n][n] denotes the natural numbers up to and including $n$, i.e., $n]=\{1,2,\dots,n\}$. Context of each arm in $\mathcal{A}$ at time $t$, $a_{t}(x)\in\mathbb{R}^{d_{a}}$, may be combined with context of user $u_{t}$ to obtain $\bm{x}_{a_{t},u_{t}}\in\mathbb{R}^{d}$, where $d=d_{u}+d_{a}$. [^6] Note that the context of the user may change across time, e.g., number of recommendations they received so far may be one of the information dimensions (Theocharous et al., , 2015). We assume that $d_{u}$ and $d_{a}$ remain unchanged throughout the time horizon. After the decision-maker (learner) pulls arm $a_{t}$ at time $t$, they receive a payoff of $r_{a_{t},u_{t}}$ from user $u_{t}$. This payoff is assumed to be a function of the context and the system state at time $t$, i.e., $r_{a_{t},u_{t}}=f(\bm{x}_{a_{t},u_{t}},\bm{S}_{t})$, where $\bm{S}_{t}$ may contain any information collected up to time $t$, and $f$ is an arbitrary function. Note that the learner has access to $\bm{S}_{0}$ prior to any interactions.
@@
Contextual bandit models provide a framework to incorporate the abundant information collected from users and items (i.e., arms) into decision making. In terms of interacting with users, by taking an action (such as recommending an item) and capturing the user’s feedback, most algorithms work in a similar manner. What distinguishes these algorithms is how they use the collected information, and how they share it among all users and actions. To materialize this, we introduce Algorithm 1 as a basis for all contextual bandit algorithms.
1Input: Set of parameters $\mathcal{C}$, set of arms $\mathcal{A}_{t}$ for $t\in[T]$, initial system information $\bm{S}_{0}$, function $g(\cdot)$
@@ 3.1 Contextual bandits in collaboration
1Input: $\alpha\in\mathbb{R}_{+}$, initial arm set $\mathcal{A}_{0}$
@@ 16  $\mathbf{b}_{a_{t}}\leftarrow\mathbf{b}_{a_{t}}+r_{a_{t},u_{t}}\bm{x}_{a_{t},u_{t}}$
1Input: $\alpha\in\mathbb{R}_{+}$, initial item set $\mathcal{A}_{0}$, user cluster set $\mathrm{C}\triangleq\cup_{i=1}^{M}\mathrm{C}_{i}$.
@@ 15 Pull arm $a_{t}=\arg\max_{a\in\mathcal{A}_{t}}p_{t,a}$ and observe reward $r_{a_{t},u_{t}}$
Notice that we associate arms with bandit parameters in Algorithm 2. To extend this to a setting where arms are more persistent than users, we instead associate users with bandit parameters and use the contextual information on arms to design the best policy. Additionally, we consider a setting where there are prohibitively many users and we remedy this by clustering users and associating these clusters with bandit parameters. In this setting, whenever a user belonging to a specific cluster gets an arm assigned to them, it is as though all the other users in the same cluster effectively do so. This also helps in propagating the information collected on a user to another one based on their perceived similarity. Assuming that a structured way to extract such similarity information exists, such as already existing friend networks in social media, collaboration propagates new information and thus speeds up the learning process, and also offers a remedy to the cold-start problem in recommender systems. The intuition is similar to that in Li et al., (2010): The payoff from a user’s interaction may be a function of both their preference and the influence of others.
There are two main approaches in the literature to inject collaboration into contextual bandit models, namely hard- and soft-clustering. In hard-clustering, connected arms are assumed to have similar or exactly the same bandit parameters. For instance, GOB.Lin (Cesa-Bianchi et al., , 2013) requires connected users in a network to share similar bandit parameters through a graph Laplacian-based regularization, i.e., the bandit parameters across connected users must be close to each other in space. CLUB (Gentile et al., , 2014) makes a similar but stronger modeling assumption and forces arms in the same cluster to have exactly the same bandit parameters. In soft-clustering, arms are connected through a weighted adjacency graph, where nodes are contextual bandits on users and the weights on the edges govern the influence between pairs of users. In upcoming parts, we introduce the notion of sparsity in these settings. We define 100% sparsity as the state where the weighted adjacency graph keeps all weights non-zero. A $50\%$ sparsity then means that the weighted adjacency graph keeps only the largest half of the weights non-zero. [^9] This is from each user’s perspective, i.e., this is a directed graph and not necessarily symmetric in both directions between nodes.
1Input: $\alpha\in\mathbb{R}_{+}$, $\mathcal{A}_{0}$, $\mathbf{W}\in\mathbb{R}^{M\times M}$.